OpenStudy (anonymous):

Solve: Square root of 128^x=32(64^x)

myininaya (myininaya):

2^7=128 2^5=32 2^6=64

OpenStudy (ybarrap):

sqrt(128^x) = 32 64^x sqrt(128^x) = 2^(6 x+5)

OpenStudy (anonymous):

Oh, sorry. The square roots only on the 128^x.

OpenStudy (ybarrap):

yes, that's what was used.